OpenStudy (anonymous):

Help? Solve using substitution. y = x - 2 x + y = 10

OpenStudy (blank ):

y=4 x=6

OpenStudy (blank ):

Tell me if you need the steps. :)

OpenStudy (anonymous):

you have y = x -2 and x + y = 10 so you plug in y into the second function. so it will be x + x -2 = 10 2x = 12 x = 6. Plug this value in any of the two function and it will give you the y-value.

OpenStudy (anonymous):

Great, thank you so much!

OpenStudy (blank ):

Who are you referring to?

OpenStudy (anonymous):

Both of you. I merely needed the answers, but the explanation will help me get the rest of the problems.